OpenStudy (anonymous):

Given the system of equations 2x-3y-9z = 21 x+3z = -12 -3x+y-4z = 21 (a) determine whether the system is inconsistent or has infinitely many solutions; Your answer is (input inconsistent or infinite) (b) if your answer is "infinite" in (a), find the complete solution. Write x and y as functions of z. x= y=

OpenStudy (anonymous):

2x - 3y - 9z = 21 -3x + y - 4z = 21--->(-3) ---------------- 2x - 3y - 9z = 21 9x - 3y + 12z = - 63 -- result of multiplying by -3 ---------------add 11x + 3z = - 42 x + 3z = -12 -->(-1) 11x + 3z = -42 -------------- -x - 3z = 12 -- result of multiplying by -1 11x + 3z = -42 -------------add 10x = - 30 x = -3 x + 3z = -12 -3 + 3z = -12 3z = -12 + 3 3z = -9 z = -3 2x - 3y - 9z = 21 2(-3) - 3y - 9(-3) = 21 -6 - 3y + 27 = 21 -3y + 21 = 21 -3y = 21 - 21 -3y = 0 y = 0 check... -3x + y - 4z = 21 x + 3z = -12 -3(-3) + 0 - 4(-3) = 21 -3 + 3(-3) = -12 9 + 0 + 12 = 21 -3 - 9 = -12 21 = 21 (correct) -12 = -12 (correct) x = -3, y = 0, z = -3 This system is consistent

OpenStudy (anonymous):

@hartnn ....can you take a peak at this...it is not inconsistent or infinite.....I think it is consistent, but that does not seem to be an answer choice.

hartnn (hartnn):

i am getting infinite solutions...

OpenStudy (anonymous):

I even checked my solutions and they turned out correct

OpenStudy (anonymous):

did I do something wrong and I am just missing it ??

hartnn (hartnn):

checking your solution now..

hartnn (hartnn):

the solution u got is one of the infinite solutions

OpenStudy (anonymous):

Oh....I just thought my solutions were the only solutions

OpenStudy (anonymous):

are you good with functions...because I am not and I cannot do the last part of the question

hartnn (hartnn):

i can try